a company has a total of 210 employees in the marketing and sales offices. for every three people employed in marketing, there are eleven people employed in sales. how many people in the office work in marketing? how many people in the office work in sales
Answer by ikleyn(53765)

.

Group the people in sets, including 3 people of marketing and 11 people of sales to each group.


The number of people in each group will be 3 + 11 = 14,

and the number of groups is  210/14 = 15.


Hence, there are 3*15 = 45 persons in marketing and 11*15 = 165 people in sales.    ANSWER

Solved.